The velocity of a bullet is reduced from 200m/s to 100m/s while travelling through a wooden block of thickness 10cm.The retardation,assuming it to be uniform will be

Grade:9

2 Answers

Aditya Garg
6 years ago
By using 2as=v^2 -u^2we haveu=200m/s v=100m/s& s=0.1m(10cm) from here deacc ( retardation)comes 150000m/s
sana
4 years ago
here initial velocity ,u = 200
final velocity , v = 100
distance, s = 10 cm = 0.1
therefore, retardation = v^2 = u^2 + 2as
substituting values, 40000 = 10000 + 0.2s
40000-10000 = 0.2s
s = 30000/0.2
s = 150000m/s^2 OR 15*10^4 m/s^2